# Spike No More: Stabilizing the Pre-training of Large Language Models

### v3

@@ Appendix G Discussion on ReLU in FFN
Appendix H Relation between Input Length and the Standard Deviation of Self-Attention
@@
We explain that the standard deviation of the self-attention layer becomes small as the input length is long. Because we assume that parameters and the input vector at each layer follow the normal distribution, the expectation of each element of $(W_{Qi}\ x)^{\mathrm{T}}(W_{Ki}\ X)$ is 00. Therefore, the expectation after the softmax function is $\frac{1}{L}$ where $L$ is the length of the input sequence. Thus, the long input sequence decreases the standard deviation of the self-attention layer.
@@
To simplify, we consider the case where the number of self-attention heads is $1$. In this case, we can obtain the variance of each calculation with the following equation.
@@
Appendix I Details on Jacobian Matrix of Self-Attention
@@
where $W_{i}\in\mathbb{R}^{d_{\mathrm{head}}\times d}$ whose corresponding element is $1$ and the others are 00. Let $J^{i}$ be the Jacobian of the $\mathrm{head}_{i}(x)$, we can represent $J^{Z}$ in Section 3.2 as follows:
@@
Based on the assumption that parameters and the vector at each layer follow the normal distribution, we assume that the mean of $D_{l}$, $E_{k}$, and $v_{l}$ is 00, and thus, we can obtain their norms from their variances. In addition, we assume that the standard deviation of $W_{Q}$, $W_{K}$ and $W_{V}$ is $\sigma$. Then, $\mathrm{var}(D_{l})=\mathrm{var}(E_{k})=dd_{\mathrm{head}}\sigma^{4}$, $\mathrm{var}(\sum_{l}^{L}D_{l})=Ldd_{\mathrm{head}}\sigma^{4}$, and $\mathrm{var}(v_{l})=d\sigma^{2}$. Thus, $\|E_{k}v_{j}^{\mathrm{T}}\|_{2}\approx\|E_{k}v_{l}^{\mathrm{T}}\|_{2}\approx\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}^{2}}$, $\|\sum_{l=1}^{L}(D_{l}v_{l}^{\mathrm{T}})\|_{2}\approx\|(\sum_{m=1}^{L}D_{m})v_{l}^{\mathrm{T}}\|_{2}\approx\sigma^{3}\sqrt{Ld^{3}d_{\mathrm{head}}^{2}}$. Therefore,
@@
Appendix J Comparison with Post-LN Transformer

### v4

@@ E.3 SwiGLU
Appendix F Relation between Input Length and the Standard Deviation of Self-Attention
@@
We explain that the standard deviation of the self-attention layer becomes small as the input length is long. Because we assume that parameters and the input vector at each layer follow the normal distribution, the expectation of each element of $(W_{Qi}\ x)^{\mathrm{T}}(W_{Ki}\ X)$ is 0. Therefore, the expectation after the softmax function is $\frac{1}{L}$ where $L$ is the length of the input sequence. Thus, the long input sequence decreases the standard deviation of the self-attention layer.
@@
To simplify, we consider the case where the number of self-attention heads is 111. In this case, we can obtain the variance of each calculation with the following equation.
@@
Appendix G Details on Jacobian Matrix of Self-Attention
@@
where $W_{i}\in\mathbb{R}^{d_{\mathrm{head}}\times d}$ whose corresponding element is 111 and the others are 0. Let $J^{i}$ be the Jacobian of the $\mathrm{head}_{i}(x)$, we can represent $J^{Z}$ in Section 3.2 as follows:
@@
Based on the assumption that parameters and the vector at each layer follow the normal distribution, we assume that the mean of $D_{l}$, $E_{k}$, and $v_{l}$ is 0, and thus, we can obtain their norms from their variances. In addition, we assume that the standard deviation of $W_{Q}$, $W_{K}$ and $W_{V}$ is $\sigma$. Then, $\mathrm{var}(D_{l})=\mathrm{var}(E_{k})=dd_{\mathrm{head}}\sigma^{4}$, $\mathrm{var}(\sum_{l}^{L}D_{l})=Ldd_{\mathrm{head}}\sigma^{4}$, and $\mathrm{var}(v_{l})=d\sigma^{2}$. Thus, $\|E_{k}v_{j}^{\mathrm{T}}\|_{2}\approx\|E_{k}v_{l}^{\mathrm{T}}\|_{2}\approx\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}^{2}}$, $\|\sum_{l=1}^{L}(D_{l}v_{l}^{\mathrm{T}})\|_{2}\approx\|(\sum_{m=1}^{L}D_{m})v_{l}^{\mathrm{T}}\|_{2}\approx\sigma^{3}\sqrt{Ld^{3}d_{\mathrm{head}}^{2}}$. Therefore,
@@
Appendix H Comparison with Post-LN Transformer